OpenStudy (anonymous):

x+3y=27 put in slope intercept form.

OpenStudy (abb0t):

slope intercept is \[\huge y=mx+b\]

OpenStudy (anonymous):

x+3y=27 put into slope intercept form.

OpenStudy (abb0t):

Are you asking me or telling me?

OpenStudy (anonymous):

I am not sure how to solve this just confused.

OpenStudy (abb0t):

rearrange it using your algebra tools. you have to get it in the form I provided: \(\large y=mx+b\) do you see how it's different from how you have it?! You have it in the form \(\large x+my=b\)

OpenStudy (anonymous):

I"m not sure if this is right y=-1x+3 Thats the problem my teacher gave me.

OpenStudy (abb0t):

Well, that's not quite right. your slope is wrong. Because remember what you do to one side, you must also do to the other. So, if you follow me this is how you do it \(\large x+my=b \rightarrow my = b-x \rightarrow y = \frac{b}{m}-\frac{x}{m}\)

OpenStudy (abb0t):

first i subtracted, then i divided...notice how i divided everything by "m"?

OpenStudy (anonymous):

Yes, so is it y=-1/3x+9

OpenStudy (anonymous):

Thanks